OpenStudy (anonymous):

if 8n+22=70, find the value of 3n.

OpenStudy (earthcitizen):

solve for n and multiply by 3

OpenStudy (anonymous):

8n+22=70 -22 -22 8n = 48 /8 /8 n = 6 3n = 3*6 = 18

OpenStudy (anonymous):

thanks that makes so much sense now esespially you surdawi. it really helps to see it written out.